OpenStudy (anonymous):

Evaluate: 10 − 64 ÷ 8 + 2^2

OpenStudy (anonymous):

24

OpenStudy (anonymous):

PEMDAS 64/8=8 10-8=2 2+22=24

jimthompson5910 (jim_thompson5910):

follow PEMDAS and go from left to right 10 − 64 ÷ 8 + 22 10 − 8 + 22 .. division 2 + 22 ... subtraction 24 ... addition

jimthompson5910 (jim_thompson5910):

so, 10 − 64 ÷ 8 + 22 = 24

jimthompson5910 (jim_thompson5910):

follow PEMDAS 10 − 64 ÷ 8 + 2^2 10 − 64 ÷ 8 + 4 ... exponentiation 10 − 8 + 4 ... division 2 + 4 ... subtraction 6 ... addition 10 − 64 ÷ 8 + 2^2 = 6